OpenStudy (anonymous):

Find the axis of symmetry: f(x) = 3(x + 4)^2 + 1 g(x) = 2x^2 –16x + 15

OpenStudy (anonymous):

what is x

OpenStudy (anonymous):

h(x)=a(x-1)-3

OpenStudy (anonymous):

Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest). those are the instructions

OpenStudy (anonymous):

where did you get h(x from?

OpenStudy (mathmale):

Find the axis of symmetry: f(x) = 3(x + 4)^2 + 1 g(x) = 2x^2 –16x + 15 Note that these are two separate parabolas, two separate questions. In f(x)=a(x-h)^2 + k, the vertex is given by (h, k). Can you use this info to find the x-coordinate of the vertex of each given parabola?

OpenStudy (anonymous):

f (x), the expression is in vertex form: a(x-h)^2 +k. The variables h and k represent the vertex so (-4, 1) The axis of symmetry is -4.

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Nice work! How about finding the ax. of sym. for the 2nd function?

OpenStudy (anonymous):

Convert it to vertex form but howww?

OpenStudy (anonymous):

@mathmale